# Task-Customized Mixture of Adapters for General Image Fusion

### v1

@@ 1 Introduction
Image fusion aims to integrate complementary information from multi-source images captured by different sensors in the same scene onto a single image. It is usually used to enhance important information and visual quality [14, 46]. Currently, general image fusion mainly includes multi-modal, multi-exposure, and multi-focus image fusion, etc. Fusion tasks exhibit diverse fusion mechanisms. The Multi-Exposure image Fusion (MEF) focuses on integrating images with multiple exposure levels into a high-quality full exposure image [38]. Each source image contributes its own illumination and structural information to the fused image [28]. The Visible-Infrared image Fusion (VIF) is a type of the Multi-Modal image Fusion (MMF) that aims at fusing the complementary information of infrared and visible modalities to produce robust and informative fused images [25]. Infrared images provide more intensity information, while visible images offer more texture and gradient information. The Multi-Focus image Fusion (MFF) aims at generating an all-in-focus image from a series of partially focused images [23]. A patch of multi-focus fused image typically corresponds to only one source image. Consequently, it can be observed that the MEF and IVF tasks involve relatively balanced fusion, while MFF is an extremely unbalanced task that tends to exhibit polarized choices.
With the rapid development of deep learning techniques, image fusion has achieved great progress in recent years [16, 30, 39], while most existing methods [28, 54, 53] focus on single image fusion scenario alone. Task-specific methods [28, 34] often employ task-specific strategies such as designing complex task-biased networks or utilizing task-specific fusion loss functions, resulting in the weak generalization to other tasks [39]. Considering that different fusion tasks have similar objectives, i.e., integrating complementary information from multiple source images, some recently proposed approaches [50, 39, 20] attempted to conduct various fusion tasks using a unified model, forming general image fusion. These methods, however, are either enmeshed in dominant-task bias [50] or multi-task commonality [39, 20] at the expense of individuality, leading to unsatisfactory performance. This motivates us to explore a more compatible fusion paradigm that is adaptively and dynamically compatible with different fusion scenarios.
@@ 2 Related Work
Mixture of Experts. Shazeer et al. [32] first proposed the mixture of experts (MoE) to increase model capacity without increasing computational complexity. Based on this, Lepikhin et al. [15] and Fedus et al. [8] integrated MoE with the transformer structure, further pushing the upper limit of network capacity. Ma et al. [24], Chen et al. [5] tackled the multi-task problem by designing multi-gates MoE. Cao et al. [3] introduces the MoE into fusion tasks, leverages foreground and background experts, to address the diversity issue of fine-grained textures for VIF fusion task. In this paper, we draw inspiration from MoE, unlike Cao et al. [3], to construct dynamic adaptive adapters for general image fusion tasks based on the visual pretrained model, resulting in our model being both sample-adaptive and task-adaptive.
@@ 4.1 Experimental Setting
Evaluation Metrics. Since the commonly used evaluation metrics for fusion tasks are not exactly the same, we customize a set of metrics for each task. Existing metrics can be mainly divided into the following four categories [38]. $\diamond$ The information theory-based metrics: entropy (EN), peak signal-to-noise ratio (PSNR), mutual information (MI), normalized MI (NMI), Nonlinear correlation information entropy ($Q_{ncie}$), feature mutual information (FMI). $\diamond$ The image feature-based metrics: phase congruency-based metric $Q_{p}$, standard deviation (SD), gradient-based metric $Q_{abf}$, and $Q_{g}$. $\diamond$ The structural similarity-based metrics: $Q_{c}$ [6], $Q_{w}$, $Q_{s}$ [36], structural similarity (SSIM), multiscale structural similarity (MS-SSIM [35]) and MEF-SSIM [27]. $\diamond$ The human perception inspired metrics: visual information fidelity ($VIF$), $Q_{cb}$ and $Q_{cv}$. Detailed information on metrics can be found in [25] and [23].
@@ 4.5 Ablation Study
We conduct ablation studies to verify the effect of our TC-MoA, MIR, and network architecture.
TC-MoA. To verify the effectiveness of TC-MoA, we remove it from our framework. As shown in the first and second rows of Table 4, the models trained by multi-task training outperform those trained for a single task, based on 5 commonly used metrics. Interestingly though, this rule reverses in the case of MEF, indicating that inter-task competition occurs for models trained directly with multiple tasks. Comparatively, the model incorporating the TC-MoA improves the performance across all metrics, suggesting that TC-MoA is dynamically compatible with different tasks.
MIR. To verify the effectiveness of MIR, we removed MIR from our framework. As shown in the last two rows of Table 4, all the evaluation metrics significantly decreased without MIR, indicating the effectiveness of MIR in learning complementary information from multi-source images. The image of the model without MIR constraint tends to exceed the fixed acceptance range of the backbone representation, resulting in a degradation of the ability to represent high-frequency information. As shown in Fig. 12, the shadow details are missed without MIR.
Adapter-based Fine-tuning. We performed ablation experiments on the backbone and adapter methods, as shown in Table 5. The fine-tuning method of “FrozenBackbone” represents freezing the entire backbone and only unfreezing the final linear layer. Our method outperforms the existing adapter-based visual fine-tuning approach AdaptFormer [4] on both base and large versions of pre-trained models, demonstrating the superiority of the TC-MoA structure in fusion tasks. More importantly,Table 9 shows our base version achieves competitive performance on the VIF task, which can be further enhanced by the large version.
@@ 4.6 Analysis and Discussion
Efficiency. We have observed the speed issues brought about by the massive parameters of the pre-trained models. To address this, we have optimized the ViT architecture by the shifted windows, which will be detailed in Sec. 8. After optimization, Table 6 shows the base and large models have accelerated by 178% and 167%, resulting in acceptable FPS (Frames Per Second) compared with other methods.
@@
Table 5: Ablation studies on fine-tuning. Base and Large represent the use of ViT-base and ViT-large structures respectively.
@@
AdaptFormer Base 0.531 0.336 0.427 0.574 0.504 0.377 0.579 0.561 0.651
TC-MoA Base 0.576 0.396 0.450 0.604 0.540 0.394 0.612 0.619 0.660
FrozenBackbone 0.330 0.238 0.363 0.452 0.408 0.333 0.450 0.428 0.596
AdaptFormer Large 0.576 0.392 0.446 0.616 0.563 0.396 0.640 0.650 0.671
TC-MoA Large 0.601 0.412 0.455 0.645 0.598 0.412 0.657 0.681 0.679
@@
Hyperparameters. We conducted hyperparameters analysis on two aspects: the number of experts ($N$) and the interval between two TC-MoAs ($\tau$). We utilize the ViT-large as the backbone, which consists of 24 transformer blocks in the encoder and 8 in the decoder. If $\tau=4$, our model contains 8 instances of TC-MoA. We employ two widely-used metrics $Q_{abf}$ and SSIM to inform the selection of hyperparameters. As presented in Fig. 8, as $N$ increases, the performance initially increases but eventually declines, with the peak performance detected at $N=4$. This suggests that increasing the number of experts does not necessarily enhance performance. Similarly, when $\tau=4$, the model demonstrates the best performance across multi-task. Hence, our experiments are conducted under $N=4$ and $\tau=4$.
@@ 6 Acknowledgement
This work was sponsored in part by the National Key R&D Program of China 2022ZD0116500, in part by the National Natural Science Foundation of China (62222608, 62106171, U23B2049, 61925602), in part by CAAI-CANN Open Fund, developed on OpenI Community, in part by the Haihe Lab of ITAI under Grant 22HHXCJC00002, in part by the Tianjin Natural Science Foundation under Grant 21JCYBJC00580, in part by Tianjin Natural Science Funds for Distinguished Young Scholar under Grant 23JCJQJC00270, and in part by the Key Laboratory of Big Data Intelligent Computing, Chongqing University of Posts and Telecommunications under Grant BDIC-2023-A-008.

### v2

@@ 1 Introduction
Image fusion aims to integrate complementary information from multi-source images captured by different sensors in the same scene onto a single image. It is usually used to enhance important information and visual quality [14]. Currently, general image fusion mainly includes multi-modal, multi-exposure, and multi-focus image fusion, etc. Fusion tasks exhibit diverse fusion mechanisms. The Multi-Exposure image Fusion (MEF) focuses on integrating images with multiple exposure levels into a high-quality full exposure image [40]. Each source image contributes its own illumination and structural information to the fused image [28]. The Visible-Infrared image Fusion (VIF) is a subfield of the Multi-Modal Fusion (MMF) [16, 8, 48] that aims at fusing complementary information of infrared and visible modalities to produce robust and informative fused images [25]. Infrared images provide more intensity information, while visible images offer more texture and gradient information. The Multi-Focus image Fusion (MFF) aims at generating an all-in-focus image from a series of partially focused images [23]. A patch of multi-focus fused image typically corresponds to only one source image. Consequently, it can be observed that the MEF and VIF tasks involve relatively balanced fusion, while MFF is an extremely unbalanced task that tends to exhibit polarized choices.
With the rapid development of deep learning techniques, image fusion has achieved great progress in recent years [31, 41], while most existing methods [28, 55, 54] focus on single image fusion scenario alone. Task-specific methods [28, 36] often employ task-specific strategies such as designing complex task-biased networks or utilizing task-specific fusion loss functions, resulting in the weak generalization to other tasks [41]. Considering that different fusion tasks have similar objectives, i.e., integrating complementary information from multiple source images, some recently proposed approaches [51, 41, 20] attempted to conduct various fusion tasks using a unified model, forming general image fusion. These methods, however, are either enmeshed in dominant-task bias [51] or multi-task commonality [41, 20] at the expense of individuality, leading to unsatisfactory performance. This motivates us to explore a more compatible fusion paradigm that is adaptively and dynamically compatible with different fusion scenarios.
@@ 2 Related Work
Mixture of Experts. Shazeer et al. [33] first proposed the mixture of experts (MoE) to increase model capacity without increasing computational complexity. Based on this, Lepikhin et al. [15] and Fedus et al. [7] integrated MoE with the transformer structure, further pushing the upper limit of network capacity. In addition, MoE has been validated as effective in other challenges. Ma et al. [24] tackled the multi-task problem by designing multi-gates MoE. Mustafa et al. [29] exploited MoE to train the contrastive learning-based vision-language foundation model. Inspired by this, we take each expert as an adapter, forming a task-customized mixture of adapters for tuning our general image fusion framework.
@@ 4.1 Experimental Setting
Evaluation Metrics. Since the commonly used evaluation metrics for fusion tasks are not exactly the same, we customize a set of metrics for each task. Existing metrics can be mainly divided into the following four categories [40]. $\diamond$ The information theory-based metrics: entropy (EN), peak signal-to-noise ratio (PSNR), mutual information (MI), normalized MI (NMI), Nonlinear correlation information entropy ($Q_{ncie}$), feature mutual information (FMI). $\diamond$ The image feature-based metrics: phase congruency-based metric $Q_{p}$, standard deviation (SD), gradient-based metric $Q_{abf}$, and $Q_{g}$. $\diamond$ The structural similarity-based metrics: $Q_{c}$ [5], $Q_{w}$, $Q_{s}$ [38], structural similarity (SSIM), multiscale structural similarity (MS-SSIM [37]) and MEF-SSIM [27]. $\diamond$ The human perception inspired metrics: visual information fidelity (VIFVIF), $Q_{cb}$ and $Q_{cv}$. Detailed information on metrics can be found in [25] and [23].
@@ 4.5 Ablation Study
We conduct ablation studies to verify the effect of our TC-MoA and network architecture.
TC-MoA. To verify the effectiveness of TC-MoA, we remove it from our framework. As shown in the first and second rows of Table 4, the models trained by multi-task training outperform those trained for a single task, based on 5 commonly used metrics. Interestingly though, this rule reverses in the case of MEF, indicating that inter-task competition occurs for models trained directly with multiple tasks. Comparatively, the model incorporating the multiple adapters improves the performance across all metrics, suggesting that TC-MoA is dynamically compatible with different tasks.
Adapter-based Fine-tuning. We performed ablation experiments on the backbone and adapter methods, as shown in Table 5. The fine-tuning method of “FrozenBackbone” represents freezing the entire backbone and only unfreezing the final linear layer. Compared to the approach of “FrozenBackbone”, our TC-MoA achieves significant fusion performance improvements by introducing a small number of learnable parameters. More importantly, our method outperforms the existing adapter-based visual fine-tuning approach AdaptFormer [4] on both base and large versions of pre-trained models, demonstrating the superiority of the TC-MoA structure in fusion tasks. In addition to this, Table 9 shows our base version achieves competitive performance on the VIF task, which can be further enhanced by the large version.
@@ 4.6 Analysis and Discussion
Efficiency. We have observed the speed issues brought about by the massive parameters of the pre-trained models. To address this, we have optimized the ViT architecture by the shifted windows, which will be detailed in Sec. 8. After optimization, for multi-task with arbitrary-size inputs, Table 6 shows the base and large versions of the model have accelerated by 178% and 167%, resulting in acceptable FPS (Frames Per Second) compared with other methods. The “Frozen” refers to the vanilla ViT architecture.
@@
Table 5: Ablation studies on fine-tuning. $\mathbb{B}$ and $\mathbb{L}$ represent the use of ViT-base and ViT-large structures respectively.
@@
FrozenBackboneB 0.175 0.089 0.275 0.250 0.139 0.227 0.295 0.189 0.466
AdaptFormerB 0.531 0.336 0.427 0.574 0.504 0.377 0.579 0.561 0.651
TC-MoAB 0.576 0.396 0.450 0.604 0.540 0.394 0.612 0.619 0.660
FrozenBackboneL 0.330 0.238 0.363 0.452 0.408 0.333 0.450 0.428 0.596
AdaptFormerL 0.576 0.392 0.446 0.616 0.563 0.396 0.640 0.650 0.671
TC-MoAL 0.601 0.412 0.455 0.645 0.598 0.412 0.657 0.681 0.679
@@
Hyperparameters. We conducted hyperparameters analysis on two aspects: the number of experts (NN) and the interval between two TC-MoAs ($\tau$). We utilize the ViT-large as our backbone, which consists of 24 transformer blocks in the encoder and 8 in the decoder. If $\tau=4$, our model contains 8 instances of TC-MoA. We employ two widely-used metrics $Q_{abf}$ and SSIM to inform the selection of hyperparameters. As presented in Fig. 8, as NN increases, the performance initially increases but eventually declines, with the peak performance detected at $N=4$. This suggests that increasing the number of experts does not necessarily enhance the model’s performance. Similarly, when $\tau=4$, the model demonstrates the best performance across multi-task. Hence, our experiments are conducted under $N=4$ and $\tau=4$.
@@ 6 Acknowledgement
This work was sponsored in part by the National Key R&D Program of China 2022ZD0116500, in part by the National Natural Science Foundation of China (62222608, 62106171, U23B2049, 61925602), in part by the Haihe Lab of ITAI under Grant 22HHXCJC00002, in part by the Tianjin Natural Science Foundation under Grant 21JCYBJC00580, in part by Tianjin Natural Science Funds for Distinguished Young Scholar under Grant 23JCJQJC00270, and in part by the Key Laboratory of Big Data Intelligent Computing, Chongqing University of Posts and Telecommunications under Grant BDIC-2023-A-008. This work was also sponsored by CAAI-CANN Open Fund, developed on OpenI Community.